# Generalizable Deepfake Detection via Effective Local-Global Feature Extraction

### v1

@@ Abstract
The rapid advancement of GANs and diffusion models has led to the generation of increasingly realistic fake images, posing significant hidden dangers and threats to society. Consequently, deepfake detection has become a pressing issue in today’s world. While some existing methods focus on forgery features from either a local or global perspective, they often overlook the complementary nature of these features. Other approaches attempt to incorporate both local and global features but rely on simplistic strategies, such as cropping, which fail to capture the intricate relationships between local features. To address these limitations, we propose a novel method that effectively combines local spatial-frequency domain features with global frequency domain information, capturing detailed and holistic forgery traces. Specifically, our method uses Discrete Wavelet Transform (DWT) and sliding windows to tile forged features and leverages attention mechanisms to extract local spatial-frequency domain information. Simultaneously, the phase component of the Fast Fourier Transform (FFT) is integrated with attention mechanisms to extract global frequency domain information, complementing the local features and ensuring the integrity of forgery detection. Comprehensive evaluations on open-world datasets generated by 34 distinct generative models demonstrate a significant improvement of 2.9% over existing state-of-the-art methods.
1 Introduction
@@
In response, various deepfake detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting Yu et al. (2020); Liu et al. (2020), to identify forgery traces. However, these approaches were often limited to training and testing within a single domain, overlooking the critical issue of cross-domain generalization. In real-world scenarios, it is essential for detection models to effectively generalize to unseen deepfake sources, including those involving unfamiliar generative models or novel categories.
Figure 1: Comparison with Previous Methods. Previous methods often extract only local information or global information to detect artifact information. However, these methods are not enough to fully capture effective artifact information. Recent methods also combine local and global information to detect forged images, but they only extract local information through simple cropping, destroying the relationship between local information. Compared with them, our method uses DWT and sliding window tiling to reasonably and effectively extract local time-frequency information, and combines the phase part of FFT to extract global frequency domain information.
In recent years, substantial efforts have been devoted to forgery detection, particularly to identifying deepfake sources that are challenging to detect in real-world scenarios, as illustrated in Figure 1 (a). To enhance generalization performance, some studies emphasize fine-grained features, such as local artifacts Haliassos et al. (2021); Chen et al. (2022) and the relationships between pixel points Chen et al. (2021); Tan et al. (2024b). Others focus on global artifact traces, including global spatial information Tan et al. (2023); Ojha et al. (2023) and frequency-domain features Tan et al. (2024a). However, prioritizing either local or global information in isolation may result in the loss of complementary details, thereby limiting generalization capabilities. To address this limitation, some methods integrate both local and global features. For example, Chen and Yang (2020) extract local features through clipping in the spatial domain while leveraging global features in the frequency space. Similarly, Wu et al. (2024) employ a local attention module and facilitates interaction with features from distant adjacent PPG graphs to capture local and global information, respectively. However, these approaches often rely on coarse image block segmentation for local feature extraction, which can lead to forced separation and loss of relevant feature relationships.
Accordingly, we propose a simple yet efficient feature extractor to capture global frequency domain and local spatial-frequency domain information, as illustrated in Figure 1 (b), which is plug-and-play and can be integrated with any CNN classifier. Specifically, our method introduces two key modules, DWT-Based Local Spatial-Frequency Domain Feature Capture Block (LoSFB) and FFT-Based Global Frequency Domain Feature Capture Block (GloFB), both using the Window Attention Block as the feature extractor. After performing the DWT, we concatenate features from four different frequency bands, each with dimensions $h\times w$. Then a $2\times 2$ sliding window is applied to extract the tiled features from each frequency band. Finally, the features within the window are tiled again, resulting in new features with dimensions $4\times(h\times w)$. At this time, when we extract the $4\times 4$ window features, they not only obtain features in four frequency domains, but also retain the local spatial-frequency information within the window, greatly improving the effect of artifact detection. Additionally, we incorporate the phase information from the FFT as a complement to the DWT, enabling the extraction of global frequency domain information. This helps guide the network to effectively learn global artifact patterns and identify more comprehensive traces of forgery. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 34 distinct models [^1] ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, SAN, AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, Stable-Diffusion v1, Stable-Diffusion v2, DALLE, Glide, Midjourney, Stable-Diffusion v1.4, Stable-Diffusion v1.5, Wukong, DALLE-2. Comprehensive experiments demonstrate that our method generalizes well across diverse and unseen sources, demonstrating a significant improvement over state-of-the-art methods by 2.9%. Our paper makes the following contributions:
  • We propose a plug-and-play feature extractor that seamlessly integrates local spatial-frequency and global frequency-domain information. This method can be effortlessly incorporated into any CNN classifier and effectively improves the model’s generalization capability.
  • We use sliding window interpretability to extract local spatial frequency information, and creatively use only phase spectrum to obtain global features to capture broader forgery artifacts.
  • Our experiments validate the effectiveness of the proposed method, which exhibits strong generalization across image detection tasks involving 34 different generation models. The method achieves a notable 2.9% improvement over state-of-the-art approaches, highlighting its superiority and robustness.
2 Related Work
In this section, we present a concise review of deepfake detection methods, mainly dividing them into three main groups: deepfake detection based on local information, deepfake detection based on global information, and deepfake detection based on local and global information fusion.
2.1 Deepfake Detection Based on Local Information
Early researches on deepfake detection often focus on specific facial regions, such as the eyes and lips, to identify forgery traces Haliassos et al. (2021); Chen et al. (2022). However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forged images. Experiments have shown that forgery traces are often embedded in fine-grained details, prompting a shift in focus toward extracting local artifact information. For instance, Chai et al. (2020) use limited receptive fields to capture local features, while Zhao et al. (2021) employ multiple spatial attention heads to enhance the network’s focus on distinct local regions. Other studies analyze forgery traces by examining relationships between local pixels. Chen et al. (2021) propose learning the similarity of local regions for a fine-grained assessment of patch similarity across scales, and Tan et al. (2024b) improve forgery detection generalization by identifying upsampling artifacts in nearby pixel grids. While capturing local artifact details, these methods overlook the fact that the overall image may still contain forged features, which are essential for generalization.
2.2 Deepfake Detection Based on Global Information
Early studies leveraged spatial information from generated images to detect fake content. For example, Yu et al. (2020); Liu et al. (2020) distinguish real and fake images using global information, such as color space and global texture. Some studies Yu et al. (2019) use GANs to generate global fingerprint information that may assist in detection. Additionally, Tan et al. (2023) and Ojha et al. (2023) extract artifact representations using global feature maps and global feature gradients, respectively. Moreover, since GAN-generated images often rely on upsampling operations that introduce artifacts at the frequency domain level, several studies have explored forgery detection in the frequency domain. For instance, Tan et al. (2024a) use FFT to extract global high-frequency information, uncovering more refined forgery features. While these methods capture sufficient global artifact information, they overlook the fact that effective artifact features often reside in subtle, localized regions. The neglect of local artifact information can significantly impair the generalization capabilities of forgery detection.
2.3 Deepfake Detection Based on Local and Global Information Fusion
Whether relying on global or local information alone, these methods detect forgery using only a single category of features, overlooking the significant information gain that arises from combining complementary features at different levels. Chen and Yang (2020) capture local features through spatial domain cropping and global features in the frequency domain. Xue et al. (2023) employ separate branches for local area feature extraction and global feature extraction, using the latter to obtain residual features. Wu et al. (2024) introduce a local attention module that interacts with distant neighboring PPG graphs to gather both local and global information. While these studies incorporate both global and local information, their approach to processing local information is limited by artificial segmentation of the face into regions, which leads to the loss of important forged details.
3 Main Backbone
The overall structure of our method is shown in Figure 2. Inspired by Mao et al. (2024), we propose an efficient local-global feature extraction framework designed to capture both global and local information. Specifically, the framework consists of two main branches: LoSFB and GloFB. Both branches utilize the Window Attention Block to extract the most relevant features, which are then fused and passed to the classifier for final classification. Notably, our approach is compatible with any classifier, allowing us to focus primarily on the feature extraction process.
Figure 2: Architecture of Our Method for Generalizable Deepfake Detection. In order to capture more comprehensive and reasonable forgery features, we use LoSFB and GloFB to capture local and global information, respectively. Specifically, we use DWT and sliding window tiling to construct local spatial-frequency features, and only use the phase part of FFT to construct global frequency domain features. Both of them use Window Attention Block to extract features. In addition, in order to weaken the information weight of global features, we add hyperparameters $\lambda$ to the global features when merging the two types of features. The final extracted features are passed into a classifier.
3.1 Window Attention Block
@@
The Window Attention Block (WAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A $1\times 1$ convolution layer is applied to extract cross-channel features, following by a $3\times 3$ depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as $X_{in}\in{\mathbb{R}}^{B\times C\times H\times W}$.
@@
LoAttention. Unlike traditional attention mechanisms, which are typically designed to capture global information, we employ a sliding window approach that divides the features into windows of size $b\times b$. By restricting the attention module to these local windows, the focus shifts to effectively extracting local features. The features are then divided into $Q,K,V\in{\mathbb{R}}^{S\times C\times N}$, where $S=B\times\frac{H}{b}\times\frac{W}{b}$ and $N=b\times b$, as required by the attention module. Since the attention operates locally within each window, the features within window $i$ are represented as $Q_{i},K_{i},V_{i}\in{\mathbb{R}}^{C\times N}$. These localized features are fed into the attention module for computation:
@@
The results of all window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
@@
After combining the outputs from LoAttention and GMLP via a dot product, we use the inverse window operation to restore the feature to its original size $B\times C\times H\times W$, followed by a convolution to map the features back to the original feature dimensions. The final output of the WAttention can be expressed as:

### v2

# Dual Frequency Branch Framework with Reconstructed Sliding Windows Attention for AI-Generated Image Detection

@@ Abstract
The rapid advancement of Generative Adversarial Networks (GANs) and diffusion models has enabled the creation of highly realistic synthetic images, presenting significant societal risks, such as misinformation and deception. As a result, detecting AI-generated images has emerged as a critical challenge. Existing researches emphasize extracting fine-grained features to enhance detector generalization, yet they often lack consideration for the importance and interdependencies of internal elements within local regions and are limited to a single frequency domain, hindering the capture of general forgery traces. To overcome the aforementioned limitations, we first utilize a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, we design a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives. Through feature enrichment of dual frequency domain branches and fine-grained feature extraction of reconstruction sliding window attention, our method achieves superior generalization detection capabilities on both GAN and diffusion model-based generative images. Evaluated on diverse datasets comprising images from 65 distinct generative models, our approach achieves a 2.13% improvement in detection accuracy over state-of-the-art methods.
Index Terms:
Deepfake detection, AI-generated image detection, AI security, Window attention.
I Introduction
@@
To prevent the abuse of generative methods, a series of AI-generated image detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting [1, 2] to identify forgery traces. However, these methods are usually limited to the domain consistent scenario, where the training and testing sets come from the same type of generative method, resulting in poor generalized detection performance in cross-generative method scenarios. With the continuous deepening of research, it is widely recognized that artifacts are often confined to subtle and localized areas. Inspired by this, extensive research has been conducted to extract generalizable and fine-grained forgery features. Specifically, Tan et al. [3] extracted more generalized artifact information from the generator’s perspective by analyzing the relationship between adjacent pixels. Other works [4, 5, 6] have further amplified the local artifacts by means of segmenting patches, thus achieving better detection performance. In addition, some studies [7, 8, 9] captured forgery traces from the perspective of frequency domain to obtain more generalized detection performance.
Although extracting local features has been widely adopted to capture generalized forgery artifacts, there are still some deficiencies as follow. 1) The existing works all lack consideration of the importance differences and interdependencies of internal elements of local regions. Although the common attention mechanism [10, 11, 12] can adaptively assign different weights to different local regions, they are not sensitive enough to the importance and interrelationship among the internal elements within the local regions. 2) Most existing works are limited to extracting local artifact features in a single frequency domain, resulting in the extracted features being insufficient for precisely capturing the general forgery traces caused by diverse generation methods.
To address above issues, it is urgent for the proposed method to tackle the following two difficulties. 1) How to model the importance difference and complex dependencies of elements within local regions? we adopt a sliding window to limit the attention mechanism to a local window and reconstruct the features within the window, making the attention mechanism more adaptively focused on pixel-level details while retaining the complex dependencies between pixels. Specifically, we concatenate features from the different frequency bands in each frequency domain. Then the nearby features are tiled in each frequency band. Finally, a sliding window is applied to extract the tiled features from all frequency bands. In this way, the reconstructed window features not only obtain features from the different frequency bands, but also retain the implicit relationships between adjacent elements. 2) How to extract more abundant and effective local features to further amplify the forgery traces? It is well known that the DWT domain can reflect local spatial frequency features containing high-frequency and low-frequency texture information [13, 14, 15]. In this paper, we find that each element of the phase part in the FFT domain contains the dependencies and structural relationship between the overall elements. That means it can capture the longer-range dependencies between the global elements, thereby enriching the extracted fine-grained local features from another perspective. To verify the rationality of above finding, we randomly selected 1k real images and 1k fake images from the dataset GenImage, and cropped them to the same size. We randomly exchange the phase part of the real image and the fake image, and use the swapped images as the input of the detector. The average prediction results of the images predicted as fake are shown in Figure 1. The smaller the prediction result is, the more likely the image is to be classified as a real image. We find that the detector tends to judge images with fake phase as fake category rather than fake amplitude. This indicates that the phase part contains more artifact traces. Moreover, the amplitude part mainly contains properties such as color and brightness, which is difficult to reflect forgery traces, as described in [16].
Therefore, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives. Based on this, a reconstructed sliding window attention mechanism is designed to obtain more fine-grained and generalized local artifact features, thus improving the detection performance across various generation methods. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 65 distinct generation models [^1] AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, SDv1, SDv2, Guided, Glide_50_27, Glide_100_10, Glide_100_27, LDM_100, LDM_200, LDM_200_cfg, DALLE, ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, ADM, Glide, Midjourney, SDv1.4, SDv1.5, VQDM, Wukong, DALLE2, R3GAN, StyleGAN3, StyleGAN-XL, StyleSwim, BlendFace, E4S, FaceSwap, InSwap, SimSwap, FLUX1-dev, Midjourney V6, GLIDE, DALLE-3, Imagen3, SD3, SDXL, BLIP, Infinite-ID, InstantID, IP-adapter, PhotoMaker, SocialRF, CommunityAI. Comprehensive experimental results show that our method achieves a significant improvement over state-of-the-art methods C2P_CLIP [17] and AIDE [18] by 5.8% and 7.0%. The overall contributions of the proposed method can be summarized as follows.
Figure 1: Average Prediction Results of the Exchange Phase. The smaller the predicted number, the more likely the model is to classify the image as real. The detector consistently judged the images with false phase as false images rather than the images with false amplitude.
  • We creatively reconstructed the sliding window attention mechanism, which effectively constructs the features within the window while limiting the attention within the local window, so that the attention mechanism can more accurately measure the importance and interdependence between internal elements in the local area.
  • We discovered that the phase part of FFT domain can reflect the artifact traces precisely in addition to the high and low frequency information of the DWT domain, and further designed the dual frequency branch framework to extract richer local forgery features from multiple perspectives.
  • Comprehensive experimental results have verified that our method exhibits strong generalization ability across 65 different generative methods. Compared with the state-of-the-art methods, the detection accuracy is improved by 2.13%, which further highlights its superiority and universality.
II Related Work
With the continuous development of generation methods, including GAN[19, 20, 21, 22], diffusion [23, 24, 25, 26], customized generation [27, 28, 29, 30, 31] and so on, the generated images are becoming more and more realistic, and the naked eye can no longer distinguish the authenticity of the images. In order to solve the future brought about by the above problems, many researchers are committed to the detection of AI-generated images. In this section, we will briefly introduce the existing detection methods from two aspects.
II-A Exploration of Fine-Grained Features in AI-Generated Image Detection
Early researches on AI-generated image detection often focused on specific facial regions, such as the eyes and lips, to identify forgery traces [32, 33]. However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forgery images. Experiments show that forgery traces are often embedded in fine details, which motivates people to turn to extracting local artifact information, including spatial and frequency domains. In terms of spatial domains, some researches narrow the focus to local areas to extract artifact information, including limited receptive fields [34, 35], cropping the image into smaller patches [4, 36, 6, 5], etc. Specifically, Chai et al. [34] employ limited receptive fields to identify patches that render images detectable, Zheng et al. [6] divides the image into multiple patches and shuffles them, and trains a block-based convolutional network for feature extraction. In addition, LGrad [8] employs pretrained CNN models as transform functions to convert images into gradients and leverages these gradients to present universal artifacts. NPR [3] improve generalization by identifying upsampling artifacts, such as pixel grid distortions, in nearby regions. In the frequency domain, F3-Net [37] introduces frequency component division and the frequency statistical difference between real images and forgery images into face forgery detection. FreDect [38] observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures. Luo et al. [39] utilize multiple high-frequency features of images to enhance generalization performance. FreqNet [9] utilize the Fast Fourier Transform (FFT) to extract global high-frequency information, revealing refined forgery features such as periodic noise or anomalies invisible in the spatial domain. AIDE [18] improved the generalization ability of forgery detection by extracting the highest and lowest fine-grained patches through DCT to capture low-level fine-grained artifact information.
Figure 2: Architecture of Our Method for Generalizable AI-Generated Image Detection. Specifically, we first design a reconstructed sliding window attention mechanism, which reconstructs the features within the sliding window and limits the attention mechanism to the local window range, forcing the attention to extract fine-grained forgery features while modeling the importance and dependencies between internal elements in the local area. At the same time, we designed a dual frequency branch framework, which facilitates the model to extract richer artifact traces from multiple perspectives by using DWT-based window tiling and the phase part of FFT. The final extracted features are passed into a classifier.
II-B Application of Attention in AI-generated Image Detection
Previous studies [10, 11, 12] has proven that: The attention mechanism can not only be trained to focus on important areas of the image, but also capture the long-range dependencies of the image, which is very useful in AI-generated image detection. Many works have explored this: Zhao et al.[40] employ multiple spatial attention heads, allowing the model to attend to distinct regions—such as the eyes and mouth—simultaneously, enhancing its ability to spot inconsistencies. Wu et al.[41] introduce a local attention module interacting with distant structures (e.g., PPG graphs), though further details from the original study would clarify this mechanism. In recent years, the attention mechanism has been widely integrated into advanced models such as Vision Transformer (ViT) [12] and CLIP [42], which not only improves the accuracy of detection, but also brings new possibilities for dealing with increasingly complex forgery techniques. For example, Ojha et al. [43] directly utilizes image features from the CLIP model for linear classification, LASTED [44] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, further advancing the field of AIGC detection, Fatformer [15] introduces a forgery-aware adapter to discern and integrate local forgery traces based on CLIP. CLIPMoLE [45] adapts a combination of shared and separate LoRAs within an MoE-based structure in deeper ViT blocks. What’s more, C2P_CLIP [17] explored the reasons and potential of CLIP’s effectiveness in AI-generated image detection, and for the first time integrated category common cues into the text encoder to inject category-related concepts into the image encoder, thereby improving detection performance.
Existing attention mechanisms, such as Vision Transformers (ViT) and CLIP, typically capture global weights and long-range dependencies to extract broad semantic information. However, they lack sensitivity to the significance and interrelationships of elements within local regions, hindering their ability to effectively detect fine-grained forgery traces. Therefore, we propose a reconstructed sliding window attention mechanism, which enhancing the feature extraction capacity of the window attention block to model the importance and dependencies of local forgery features. In addition, a dual frequency branch network is designed to enrich the features within the window from different angles, which promotes the extraction of more comprehensive local forgery features in the window.
III The Proposed Method
III-A Main Backbone
The overall structure of our method is shown in Figure 2. Specifically, we first propose a reconstructed sliding window attention mechanism, which reconstructs the features within the window while limiting the attention to the local window, forcing the attention to extract fine-grained forgery traces while modeling the importance and interdependence between internal elements within the local range. In addition, in order to enrich the extraction of local features from different perspectives, we design a dual frequency branch framework, using DWT-based window tiling and phase part of FFT before using window attention. In this way, the window feature not only obtains the features of different frequency bands, but also retains the implicit relationship between adjacent elements, while complementing the long-range dependency between features, which improves the overall generalization ability of the model.
Figure 3: After DWT, sliding window tiling is used to tile the features of each channel into a shape of $4\times(h\times w)$. In the subsequent LoAttention module, $\times 4$ 4 features are extracted as shown in the blue box, which not only obtain features from the different frequency bands, but also retain the implicit relationships between adjacent elements.
III-B Reconstructed Sliding Window Attention Block
@@
The Reconstructed Sliding Window Attention Block (RSWAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A $\times 1$ 1 convolution layer is applied to extract cross-channel features, following by a $\times 3$ 3 depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as $X_{in}\in{\mathbb{R}}^{B\times C\times H\times W}$.
@@
LoAttention. The attention mechanism obtains better results by training different weights of features. However, the weights of traditional attention mechanisms are derived from the features of the entire image, which contains a lot of semantic information. This weight design may cause the model to be disturbed when extracting fine-grained forgery traces. Thus we designed LoAttention module. Specifically, we employ a sliding window approach that divides the features into windows of size $b\times b$. By restricting the attention module to these local windows, the focus shifts to effectively extracting local features. The features are then divided into $Q,K,V\in{\mathbb{R}}^{S\times C\times N}$, where $S=B\times\frac{H}{b}\times\frac{W}{b}$ and $N=b\times b$, as required by the attention module. Since the attention operates locally within each window, the features within window $i$ are represented as $Q_{i},K_{i},V_{i}\in{\mathbb{R}}^{C\times N}$. These localized features are fed into the attention module for computation:
@@
The results of all sliding window attention operations can be expressed as $A=\{A_{1},A_{2},\cdots,A_{S}\}\in{\mathbb{R}}^{S\times C\times N}$.
@@
After combining the outputs from LoAttention and GMLP via a dot product, we use the inverse window operation to restore the feature to its original size $B\times C\times H\times W$, followed by a convolution to map the features back to the original feature dimensions. The final output of the RSWAttention can be expressed as: